OpenStudy (anonymous):

A triangle has sides of lengths 18, 80, and 82. is ii a right triangle? Explain answer choices: a. yes, 18^2+80^2=82^2 b.yes, 18^2+80^2=82^2 c.no, 18^2+80^2=82^2 d.no,18^2+80^2=82^2(pretend that the equal sign is crossed out)

OpenStudy (anonymous):

@elementwielder @Abbles @McDonaldsGirl<3 @TheRealMeeeee

OpenStudy (anonymous):

are a and b the same?

OpenStudy (abbles):

You need to use the Pythagorean theorem to figure this out. a^2 + b^2 = c^2

OpenStudy (abbles):

18^2 + 80 ^2 = 82^2 Simplify this and check. If it doesn't work out, then the triangle is not a right triangle.

OpenStudy (anonymous):

a and b are not the same one has a normal = sign the other should be crossed out

OpenStudy (anonymous):

18^2+80^2=82^2 is true

OpenStudy (anonymous):

@iambatman

OpenStudy (abbles):

What's the difference between A and B? It's one of those, the one with the regular = sign.

OpenStudy (anonymous):

ya i was wondering that earlier

OpenStudy (anonymous):

a should be a =with a/ over i b is a regular = sign

OpenStudy (abbles):

The answer is B then.

OpenStudy (anonymous):

then b is the answer

OpenStudy (anonymous):

should be a = sign with a / over it*

OpenStudy (anonymous):

oh ok then thanks @Abbles and @elementwielder

OpenStudy (abbles):

yep :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

gn u guys

OpenStudy (anonymous):

nite